# Exposía: Academic Writing Assessment of Exposés and Peer Feedback

### v1

@@ Abstract
We present Exposía, the first public dataset [^1] https://tudatalib.ulb.tu-darmstadt.de/handle/tudatalib/4978 that connects writing and feedback assessment in higher education, enabling research on educationally grounded approaches to academic writing evaluation. Exposía includes student research project proposals and peer and instructor feedback consisting of comments and free-text reviews. The dataset was collected in the “Introduction to Scientific Work” course of the Computer Science undergraduate program that focuses on teaching academic writing skills and providing peer feedback on academic writing. Exposía reflects the multi-stage nature of the academic writing process that includes drafting, providing and receiving feedback, and revising the writing based on the feedback received. Both the project proposals and peer feedback are accompanied by human assessment scores based on a fine-grained, pedagogically-grounded schema for writing and feedback assessment that we develop.
We use Exposía to benchmark state-of-the-art open-source large language models (LLMs) for two tasks: automated scoring of (1) the proposals and (2) the student reviews. The strongest LLMs attain high agreement on scoring aspects that require little domain knowledge but degrade on dimensions evaluating content, in line with human agreement values. We find that LLMs align better with the human instructors giving high scores. Finally, we establish that a prompting strategy that scores multiple aspects of the writing together is the most effective, an important finding for classroom deployment.
Exposía: Academic Writing Assessment of Exposés and Peer Feedback
@@ 1 Introduction
Figure 1: Overview of Exposía. In the university course “Introduction to Scientific Work”, students submit a draft exposé, receive feedback in the form of comments and a free-text review, revise and submit a final exposé. Feedback is produced by peers and instructors. Instructors grade draft and final exposés and student feedback.
@@
Writing a research project proposal [^2] We will use research project proposal, research project exposé, or exposé interchangeably in the paper. is the number one challenge in academic education and represents a pressing concern across the academic community Muneer et al. (2020). It is also an essential skill required for conducting research and for achieving academic success at all levels Kivunja (2016); Grech (2018). However, a lot of students lack the necessary competencies to develop a research proposal and its core components, including identifying a research question, formulating appropriate research methodology, and defining hypotheses and methods to solve the problem Rastri et al. (2023).
@@
At the same time, higher education and research in general face another pressing challenge: the peer review crisis Lu et al. (2025). High-quality feedback is an essential pedagogical tool that has been shown to vastly improve learning outcomes of both authors and reviewers Wu and Schunn (2021). Yet, students are rarely taught this skill in a classroom setting. Together, these challenges highlight the need for instructional approaches that simultaneously support research project proposal writing and developing effective peer review skills.
Teaching students how to write an effective project exposé and how to provide high-quality feedback is crucial in university education for several reasons. First, assessing both project exposés and peer feedback requires substantial human effort and does not scale well Grabau and Wilson (1995). Second, developing automated assessment tools requires modeling complex cross-document relationships, connecting student writing, feedback, and revisions. Moreover, peer feedback is often of low quality and highly inconsistent Purkayastha et al. (2025). Finally, collecting data in an educational setting is extremely challenging (Section 3). As a result, existing resources for computational writing assessment are fragmented across subtasks such as automated essay scoring Ke and Ng (2019), peer feedback analysis Staudinger et al. (2024), or revision modeling Ruan et al. (2024).
@@
In this work, we present Exposía, the first corpus that integrates drafting, feedback, revision, and fine-grained exposé and feedback assessment scores within a single dataset (Figure 1). Exposía is continuously collected in a newly developed university Computer Science course. The course focuses on teaching students (1) how to write a research project exposé and (2) how to write peer feedback. While this is an undergraduate course, the skills it develops are equally applicable to graduate students and other early-career researchers.
@@
To the best of our knowledge, Exposía is the first dataset, collected in an educational setting, that features this unique class of cross-document relations, enabling new applications for writing and grading assistance. Because the corpus contains paired drafts and final exposés, it enables quantifying draft-to-final improvements on specific criteria and relating these changes to the feedback received. Overall, we find that exposés improve in response to feedback, particularly on criteria related to the in-domain content of the writing rather than form. We further establish baseline models with open-source LLMs for two tasks: exposé scoring and feedback scoring, based on the fine-grained grading schema for writing and feedback assessment.
@@
The main contribution of this work is the introduction of a novel problem definition of exposé and review scoring in the educational context. The result is the first public dataset of student research exposés and feedback, accompanied by human assessment scores on exposés and feedback, based on the fine-grained grading schema that we develop. We present in-depth analyses of the dataset, and models for the corresponding tasks of exposé and feedback scoring, establishing baselines that the community can build on.
@@ 2 Related Work
Automated Essay Scoring
The task of Automated Essay Scoring (AES) Page (1966), aims to reduce human grading effort and improve consistency in writing assessment. Most widely used AES benchmarks consist of essays written to standardized prompts in general-domain language. These essays are typically manually annotated with a holistic score, complemented by scores for a small set of assessment criteria (e.g., content, organization, language) Mathias and Bhattacharyya (2018); Gaudeau (2025); Li and Ng (2024); Xue et al. (2021); Crossley et al. (2023); Ishikawa (2023). Importantly, these benchmarks do not include human feedback, revisions, or feedback assessment scores.
[…]
LLM-based scoring and transfer to peer review Recent work evaluates LLMs as scorers and emphasizes that educational use requires evidence beyond accuracy Pack et al. (2024); Huang and Wilson (2025). Exposía complements this line of work, by enabling fine-grained exposé and feedback assessment with LLMs and interpreting LLM–human alignment relative to human agreement (Section 6). Because our corpus links each exposé draft with feedback text and assessment scores, it is related to scholarly peer-review datasets Kang et al. (2018); Dycke et al. (2023). However, these datasets provide only coarse-grained annotations. Importantly, they serve fundamentally different purposes, as they focus on scientific communication in a professional decision-making context, whereas Exposía is situated in an educational environment with a focus on learning and assessment. This creates a testbed for transfer experiments between educational feedback assessment and scholarly reviewing.
@@ 3 Course Setting and Study Protocol
Data collection protocol
The dataset was collected in the newly-introduced Bachelor-level Computer Science course Introduction to Scientific Work offered for the first time in winter term 2024. The corpus is the result of a targeted data collection. [^3] The institutional IRB has approved the related proposal. Collecting data in an educational setting is challenging due to the high sensitivity of the data: collection, storage, and sharing of educational data raise serious privacy and ethical issues Guan et al. (2023).
@@
In the beginning of the course, [^5] The syllabus is provided in Appendix A. alongside the lectures, the students received guidelines for writing an exposé (Appendix B) and the exposé grading criteria (Section 4.2). Each student produced an exposé draft. This was followed by a peer-review phase, where each student wrote double-blind reviews for two exposés written by their peers, selected uniformly at random. The students were provided with review grading criteria (Section 4.3). The instructional staff also provided a review for each exposé. This was followed by a revision phase to produce the final exposé, based on the feedback. Both the drafts and final versions, and the student reviews were graded by the instructional staff. The final grade was calculated as follows: draft exposé 30%, peer reviews 30%, and final exposé 40%. [^6] Grades for each component are derived from criterion-based scores (see Section 4 and Appendix F.3).
Reviewing was facilitated through the open-access CARE platform Zyska et al. (2023, 2025).
Continued large-scale data collection 193 students enrolled in the course, [^7] Enrollment was limited because the course was offered for the first time as a pilot. of whom 159 completed all requirements. The Exposía dataset refers to the open data that participants agreed to contribute for public release. All 12 instructors and about 30% of students consented to share their materials.
Exposía currently comprises 55 exposés, 306 reviews, and 2,253 inline comments.
We emphasize that the dataset will continue to grow, as the course will be offered annually. We will continue to expand and update the dataset with additional data. In the winter term 2025, 600 students enrolled in the course. We expect at least 30% of the students to consent to the release of their materials, consistent with the previous semester.
In general, 600 students are expected to enroll each year, with a higher proportion of students willing to contribute their content, as they understand the significance of data collection in the course that teaches important skills for writing scientific papers and for providing effective feedback on scientific writing. This is consistent with similar important collection efforts (Dycke et al. (2023)).
@@ 4.1 Exposía Dataset Overview
Figure 2 illustrates two main components of the Exposía: exposés and feedback.
@@ 4.2 Exposé Assessment Criteria and Rubrics
Both the draft and the final exposé are graded independently by an instructional staff member on 36 criteria and include 36 per-criterion assessment scores (Appendix Table LABEL:tab:expose_rubric). Each criterion is rated on an ordinal scale of [0,1], [0,1,2], or [0,-1,-2].
Designing exposé assessment criteria and rubrics The extensive criterion-based grading scheme is designed to support exposé assessment across multiple aspects. These criteria were derived using 18 guidance documents for writing thesis exposés from major universities in Germany and Austria (Appendix E). These were supplemented with over 100 exposés previously submitted in the Computer Science Department. The 36 criteria center on the different aspects of writing a scientific proposal, with a focus on content assessment.
Guided by research in rubric construction and fairness-oriented assessment Jonsson and Svingby (2007); Mayring (2000); Kuckartz and Rädiker (2023), we organized the criteria into nine high-level categories referred to as rubrics (Table 1) that decompose evaluation into orthogonal dimensions, by grouping criteria that evaluate related aspects of the exposé writing.
Rubrics are an effective assessment tool and are constructed to ensure consistency in evaluation across a diverse student body and multiple evaluators Northern Illinois University Center for Innovative Teaching and Learning (2012). See Appendix D.1 for more details.
Code Rubric Description #Crit.
Exposé
LA Lang./Quality Assessment of language quality and structural coherence of the text. 2
ME Metadata Evaluation of title appropriateness and creativity. 2
ST Form/Structure Evaluation of document formatting and structural requirements. 2
MO Motivation Assessment of problem identification, context establishment, and research question formulation. 7
AP Approach Assessment of state of the art analysis, theoretical framework. 8
MD Methodology Assessment of the methodology. 7
SC Schedule Assessment of project timeline structure, completeness, and realism. 4
BI Bibliography Assessment of bibliography consistency and literature relevance. 2
AD Addt’l points Add points for great submissions or deduct points for major mistakes. 2
Review
SC Structure/Clarity Assessment of overall structure and clarity. 2
LG Language Assessment of language quality and tone. 2
FU Feed Up Assessment of learning goals and success criteria. 2
FB Feed Back Assessment of feedback quality and knowledge transfer. 6
$F$ Feed Forward Assessment of future learning guidance. 2
CQ Content Quality Assessment of content accuracy and usefulness. 3
ER Errors Assessment of potential errors and issues in feedback. 5
AD Addt’l points Add points for great reviews or deduct points for major mistakes. 2
Table 1: Exposé and review grading rubrics. The last column shows the number of criteria in the rubric.
@@ 4.3 Review Assessment Criteria and Rubrics
Each student review was independently graded by an instructional staff member and received 24 assessment scores, in line with the 24-criterion grading scheme (Appendix Table LABEL:tab:review_rubric) organized into eight rubrics (Table 1), with each criterion rated on an ordinal scale of [0,1], [0,1,2], or [-1,1]. The rubrics are shown in Table 1.
Designing review assessment criteria and rubrics The main goal of assessing reviews was to capture the extent to which peer feedback supports learning, reflection, and revision. Accordingly, the rubrics operationalize key educational principles of effective feedback—clarity, constructiveness, specificity, and usefulness—through an analytic, criterion-based framework that is transparent, fair, and aligned with established feedback theory Hattie and Timperley (2007). As such, our rubric design is primarily informed by the feedback model in Hattie and Timperley (2007), which conceptualizes feedback as answering three questions: Where am I going? (Feed Up), How am I going? (Feed Back), and Where to next? (Feed Forward).
Drawing on Narciss (2008), we integrated the components of feedback as criteria within the 8 rubrics. We further incorporated insights from D’Arcy et al. (2024), which analyzes revisions in response to peer reviews in scientific writing, highlighting that actionable comments, those offering concrete, suggestions, are more likely to lead to substantive text revisions. As such, our rubrics emphasize argumentation, accuracy, and clear suggestions for improvement as indicators of pedagogically effective feedback Shute (2008).
Finally, while actionable feedback promotes learning, ineffective or misleading reviews can have the opposite effect. To this end, inspired by Du et al. (2024), the criteria include a penalty for unhelpful feedback patterns (e.g., neglect of key elements, unsupported claims, or contradictory advice). Appendix D.2 provides more detail.
@@ 5.1 Exposía Dataset Statistics
The dataset statistics are in Table 2. As the dataset comprises only materials for which consent was granted, some exposés lack the full set of three reviews (Table 18), and there exist 184 orphaned reviews because the corresponding exposés could not be made public. The total number of student reviews is 110, including 34 reviews with an existing exposé and 76 that are orphaned. Each draft exposé includes on average 41 inline comments. Weakness is the most frequent tag (59%) of all comments, followed by Strength (18%), Other (14%) and Highlight (9%).
@@
Figure 4: Human–LLM agreement (QWA) for exposé scoring by expertise level. The dotted line shows human-human QWA.
Figure 5: Group asymmetry in human–LLM agreement by exposé rubric. Each point shows the difference in agreement between an LLM and human raters (Group 1 (G1) vs. Group 2 (G2)). The vertical line at $\Delta=0$ indicates equal agreement of the LLM with human G1 and G2 raters. Points to the right indicate the model agrees more with G1 raters than with G2 raters.
@@ 5.2 Inter-Annotator Agreement
Table 3: IAA on exposé and review scoring. Quadratic weighted agreement (QWA), Krippendorff’s ($\boldsymbol{\alpha}$), and Pearson $r$. $r$ for Addt’l points (review) is undefined due to zero variance; we exclude it from the average. Result for each rubric is averaged over all criteria in the rubric. The number of criteria per rubric is next to rubric name.
We compute inter-annotator agreement (IAA) on exposé and review criterion-based scores. Each draft exposé and student review was independently graded by two instructors on 36 exposé and 24 review scoring criteria, respectively.
Rater training During the semester, each draft exposé and student review was graded by an instructor only once; after the course, each item was graded again for calculating IAA. We refer to instructors who graded during the semester as Group 1. They attended weekly sessions and discussed edge cases. Group 1 raters were instructed to resolve uncertainty in favor of the student, because grades were at stake. Group 2 includes a different subset of instructors (with one who also participated in Group 1). They received the same guidelines, but they were not instructed to resolve uncertainty in favor of the student. See Appendix G.1 for more details.
IAA metrics We report Krippendorff’s $\alpha$ Krippendorff (2019), [^8] Krippendorff’s $\alpha$ can be used in a setup such as ours, where we are comparing multiple raters where raters also belong to one of the two groups. Pearson correlation coefficient $r$, and raw quadratic weighted percent agreement (QWA). We have several criteria with skewed distributions where the majority label accounts for over 90% cases (Tables 19 and  20). In this case, both $\alpha$ and $r$ can be misleading Jeni et al. (2013); Zhao et al. (2022), and QWA is the most informative metric. Appendix G.2 and G.3 provide more detail.

### v2

# Exposía: Teaching and Assessment of Academic Writing Skills for Research Project Proposals and Peer Feedback

@@ Abstract
We present Exposía, the first public dataset [^1] https://tudatalib.ulb.tu-darmstadt.de/handle/tudatalib/4978 that connects writing and feedback in higher education, enabling research on educationally grounded computational approaches to teaching and evaluating academic writing. Exposía includes student research project proposals and peer and instructor feedback consisting of comments and free-text reviews. The dataset was collected in the “Introduction to Scientific Work” course of the Computer Science program that focuses on teaching academic writing skills and providing peer feedback on academic writing. Exposía reflects the multi-stage nature of the academic writing process that includes drafting, receiving feedback, and revising the writing based on the feedback received. Both the project proposals and peer feedback are accompanied by human assessment scores based on a fine-grained, pedagogically-grounded schema for writing and feedback assessment that we develop.
We use Exposía to benchmark state-of-the-art large language models (LLMs) on two tasks: automated scoring of (1) the proposals and (2) the student reviews. We find that the two tasks benefit from different LLMs. Furthermore, closed-source models consistently outperform open-weight models, motivating further research on improving the performance of open-weight models preferred in classroom settings. Finally, we establish that a prompting strategy that scores multiple aspects of the writing together is the most effective, an important finding for classroom deployment.
Exposía: Teaching and Assessment of Academic Writing Skills for Research Project Proposals and Peer Feedback
@@ 1 Introduction
Figure 1: Overview of Exposía. Students submit a draft exposé, receive feedback from instructors and peers in the form of comments and free-text reviews, revise and submit a final exposé. Instructors grade draft and final exposés and student feedback.
@@
Writing a research project proposal or exposé [^2] The term exposé is used in many universities to refer to a research project proposal, a structured document that outlines a planned research project. We will use research project proposal, (research project) exposé and exposé interchangeably. is the number one challenge in academic education and represents a pressing concern across the academic community Muneer et al. (2020). It is also an essential skill required for conducting research and for achieving academic success Kivunja (2016); Grech (2018). However, many students lack the necessary competencies to develop a research proposal and its core components, including identifying a research question, formulating appropriate research methodology, and defining hypotheses and methods to solve the problem Rastri et al. (2023).
@@
At the same time, higher education faces another pressing challenge: the peer review crisis Lu et al. (2025). Peer feedback is a powerful pedagogical tool that improves the writing performance of both authors and reviewers Wu and Schunn (2021); Jiang et al. (2024). Yet, students are rarely taught this skill in a classroom. Notably, research shows a strong link between academic writing skills and the ability to provide effective feedback Huisman et al. (2019), underscoring the need for instructional approaches that simultaneously support research project proposal writing and the development of effective peer review skills.
Teaching students how to write a good exposé and how to provide high-quality feedback is very challenging. Formative assessment, which is used during the learning process, is a fundamental component of effective teaching Black and Wiliam (1998); Gass and Mackey (2012). However, human evaluation of exposés and peer feedback requires substantial effort Grabau and Wilson (1995). This applies to holistic scoring of exposés and feedback, fine-grained aspect-based assessment, and especially to the provision of textual feedback in the form of comments and reviews. These limitations necessitate the development of computational approaches – not only for automated scoring of writing and feedback quality, but also for generating detailed text-level feedback on academic writing that evaluates and supports learning. In addition to the lack of appropriate benchmarks both for holistic and fine-grained scoring, providing textual feedback requires modeling complex cross-document relationships between proposal drafts, feedback, and resulting revisions. Curating such data is a time-consuming process. Finally, collecting data in a live course comes with ethical and legal challenges (see Limitations). As a result, existing resources for computational writing assistance and assessment are fragmented across subtasks such as automated essay scoring Ke and Ng (2019), peer feedback analysis Staudinger et al. (2024), or revision modeling Ruan et al. (2024).
@@
We present Exposía, the first corpus that integrates drafting, feedback, revision, and fine-grained exposé and feedback assessment scores within a single dataset (Figure 1). Exposía is continuously collected in a newly developed university Computer Science course. The course focuses on teaching students (1) how to write a research project exposé and (2) how to write peer feedback. While this is an undergraduate Computer Science course, the skills it develops generalize across disciplines (Section 4) and are equally applicable to graduate students and other early-career researchers.
@@
To the best of our knowledge, Exposía is the first dataset, collected in an educational setting, that features this unique class of cross-document relations, enabling new applications and tasks for teaching the skills for writing research proposals and providing effective feedback (see Section 7). We further establish baseline models with LLMs for two tasks in this space: exposé scoring and feedback scoring, based on the fine-grained grading schema for writing and feedback assessment.
@@
The main contribution of this work is the introduction of a novel problem definition for teaching and assessment of exposé and feedback writing skills in an educational setting. The result is the first unique public dataset of student research exposés and feedback, both accompanied by human assessment scores, based on the fine-grained grading schema that we develop. We present in-depth analyses of the dataset, and models for the corresponding tasks of exposé and feedback scoring, establishing baselines that the community can build on.
@@ 2 Related Work
We focus on exposé and feedback writing skills, which are distinct from other (academic) writing skills. Below, we discuss related research directions and the gaps in existing datasets.
Automated essay scoring (AES) AES Page (1966) is an established NLP task that focuses on writing assessment. However, AES benchmarks consist of persuasive essays in general-domain language, manually annotated with a holistic score, complemented by scores for a small set of assessment criteria (e.g., content, organization, language) Mathias and Bhattacharyya (2018); Gaudeau (2025); Li and Ng (2024); Xue et al. (2021); Crossley et al. (2023); Ishikawa (2023). Importantly, these benchmarks do not include human feedback, revisions, or feedback assessment scores.
Peer feedback Because Exposía links each exposé draft with feedback text and assessment scores, it is related to scholarly peer-review datasets Kang et al. (2018); Dycke et al. (2023). However, these datasets provide only coarse-grained annotations. Crucially, they serve fundamentally different purposes, as they focus on scientific communication in a professional decision-making context, whereas Exposía is situated in an educational environment with a focus on learning and assessment.
Connecting writing, feedback, and revisions A small number of resources connect student writing with peer or instructor feedback and/or revision Zhang et al. (2017); Kashefi et al. (2022); Pilan et al. (2020); Rietsche et al. (2022), enabling analyses of how feedback relates to revision outcomes. However, existing datasets still focus on persuasive essays in general-domain language. Furthermore, they lack fine-grained annotations that allow for evaluating the impact of feedback on revisions and writing outcomes, and that are provided in Exposía: (i) passage-anchored inline comments tied to draft spans, (ii) role-separated peer and instructor feedback, and (iii) criterion-level assessment scores on the text to be assessed and the feedback.
@@ 3 Course Setting and Study Protocol
The dataset was collected in the newly-introduced Bachelor-level Computer Science course Introduction to Scientific Work in winter 2024.333The institutional IRB has approved the related proposal.
@@
In the beginning of the course, [^5] The syllabus is provided in Appendix A. alongside the lectures, the students received guidelines for writing an exposé (Appendix B) and the exposé grading criteria (Section 4.2). Each student produced an exposé draft. This was followed by a peer-review phase, where each student wrote double-blind reviews for two exposés written by their peers, selected uniformly at random. The students were provided with review grading criteria (Section 4.3). The instructional staff also provided a review for each exposé. This was followed by a revision phase to produce the final exposé, based on the feedback. Both the drafts and final versions, and the student reviews were graded by the instructional staff. The final grade was calculated as follows: draft exposé 30%, peer reviews 30%, and final exposé 40%. [^6] Grades for each component are derived from criterion-based scores (see Section 4 and Appendix F.3). Reviewing was facilitated through the open-access CARE platform Zyska et al. (2023, 2025).
Out of 193 students enrolled in the course, [^7] Enrollment was limited because the course was offered for the first time as a pilot. 159 completed all requirements. The Exposía dataset refers to the open data that participants agreed to contribute for public release. All 12 instructors and 30% of students consented to share their materials.
Continued large-scale data collection Exposía currently comprises 55 exposés, 306 reviews, and 2,253 inline comments. The course will run annually, and the dataset will be expanded with new data. In the winter term 2025, about 500 students enrolled in the course, and 30% of the students consented to the release of their materials. This ongoing effort aligns with similar important collection efforts Dycke et al. (2023). See also Limitations.
@@ 4.1 Exposía Dataset Overview
Figure 2 illustrates two main components of the Exposía dataset: exposés and feedback.
@@
Grading of the exposés and student reviews The exposés and student reviews are graded by instructors on 36 and 24 assessment criteria, respectively. These criteria are designed to reflect traits that could impact an overall quality of the exposé and review, respectively, and were developed in consultation with the instructors of the course. The criteria are grounded in established research on exposé and feedback writing, as detailed below (see Appendix D for annotation guidelines). Before the course was run, the criteria were assessed by the instructors and their definitions were adjusted for clarity (for more detail, see Appendix G.1).
@@ 4.2 Exposé Assessment Criteria and Rubrics
Both the draft and the final exposé are graded independently by an instructional staff member and include 36 per-criterion assessment scores (see Table LABEL:tab:expose_rubric for guidelines on grading). Each criterion is rated on an ordinal scale: [0,1], [0,1,2], or [0,-1,-2].
Designing exposé assessment criteria and rubrics The criteria were derived using 18 cross-disciplinary guidance documents for writing thesis exposés from major universities in Germany and Austria (Appendix E), supplemented by observations from instructors in our institution, based on over 100 exposés previously submitted in the Computer Science Department. The 36 criteria center on the different aspects of writing a scientific proposal, with a focus on content assessment. Importantly, the criteria are topic-independent (rather than specific to the research topic of the proposal). This approach is in line with aspect-based assessment used in automated essay scoring  Li and Ng (2024) and focuses assessment on transferable proposal-writing skills (e.g., motivation, methodology, etc.). Accordingly, the same 36 criteria are applied uniformly across the proposal topics.
Guided by prior research in rubric construction and fairness-oriented assessment Jonsson and Svingby (2007); Mayring (2000); Kuckartz and Rädiker (2023), we organize the criteria into nine high-level categories referred to as rubrics [^8] Rubrics are an effective assessment tool and are constructed to ensure consistency in evaluation across a diverse student body and multiple evaluators Northern Illinois University Center for Innovative Teaching and Learning (2012). See Appendix D.1 for more details. (Table 2), by grouping together criteria that assess related aspects of the exposé writing.
@@ 4.3 Review Assessment Criteria and Rubrics
Each student review was independently graded by an instructional staff member on 24 criteria (see Table LABEL:tab:review_rubric for the guidelines on grading) organized into eight rubrics (Table 2), with each criterion rated on an ordinal scale of [0,1], [0,1,2], or [-1,1].
Designing review assessment criteria and rubrics The review assessment criteria are grounded in established pedagogical theory and, similar to the exposé criteria, focus on feedback skills that are not tied to the proposal topic (e.g., argumentation, actionable suggestions). The criteria operationalize key educational principles of effective feedback and are primarily informed by the feedback model in Hattie and Timperley (2007).
Drawing on Narciss (2008), we integrate the components of feedback as criteria within the eight rubrics. We further incorporate insights from D’Arcy et al. (2024), which analyzes revisions in response to peer reviews in scientific writing, highlighting that actionable comments, those offering concrete, suggestions, are more likely to lead to substantive text revisions. As such, our rubrics emphasize argumentation, accuracy, and clear suggestions for improvement as indicators of pedagogically effective feedback Shute (2008). Appendix D.2 provides more detail.
@@ 5.1 Exposía Dataset Statistics
The dataset statistics are in Table 1. As the dataset comprises only materials for which consent was granted, some exposés lack the full set of three reviews (Table 16), and there exist 184 orphaned reviews because the corresponding exposés could not be made public. The number of student reviews is 110, including 76 that are orphaned. Each draft exposé includes on average 41 inline comments. Weakness is the most frequent tag (59%) of all comments, followed by Strength (18%), Other (14%) and Highlight (9%).
@@ 5.2 Inter-Annotator Agreement
Table 2: IAA on exposé and review grading. Quadratic weighted agreement (QWA), Krippendorff’s ($\boldsymbol{\alpha}$), and Pearson $r$. $r$ for Addt’l points (review) is undefined due to zero variance and excluded from the average. Result for each rubric is averaged over all criteria in the rubric. The number of criteria per rubric is next to rubric name. Rubric definitions are in Tables LABEL:tab:expose_rubric and LABEL:tab:review_rubric.
We compute inter-annotator agreement (IAA) on exposé and review criterion-based scores: each draft exposé and student review are independently graded by two instructors on 36 exposé and 24 review criteria, respectively (using two raters to compute IAA is common practice in similar settings such as AES Li and Ng (2024)), yielding 3,960 exposé assessment scores and 6,720 review assessment scores. These numbers ensure statistical robustness for reliability analysis.
Rater training and annotation effort During the semester, each draft exposé and student review were graded by an instructor only once; after the course, each item was graded again for calculating IAA. This annotation effort is substantial: grading one exposé requires assessing it on 36 criteria and providing feedback (inline comments and a review), taking 1–2 hours in total. Grading one review involves assessing it on 24 criteria and takes 20–40 minutes (see Appendix F.3 for more detail). See Appendix G.1 for more detail on rater training.
IAA metrics We report Krippendorff’s $\alpha$ Krippendorff (2019), [^9] Krippendorff’s $\alpha$ can be used in a setup such as ours, where we are comparing multiple raters where raters also belong to one of the two groups. Pearson correlation coefficient $r$, and raw quadratic weighted percent agreement (QWA). We have several criteria with skewed distributions where the majority label accounts for over 90% cases (Tables 18 and  19). In this case, both $\alpha$ and $r$ can be misleading Jeni et al. (2013); Zhao et al. (2022), and QWA is the most informative metric. Appendix G.2 and G.3 provide more detail.